OpenStudy (anonymous):

Solve. t divided by 2 is less than or equal to 4 t is less than or equal to 2 t is less than or equal to 4 t is less than or equal to 6 t is less than or equal to 8

OpenStudy (nincompoop):

if you have 1/2 of something, what do you do to make it whole or 1?

OpenStudy (nincompoop):

I guess you're just here looking for an answer ...

OpenStudy (anonymous):

u add the same amount

OpenStudy (mathmale):

Let's look at question #1 again: "t divided by 2 is less than or equal to 4". Write this in mathematical symbols:\[\frac{ t }{ 2 }\le4\]

OpenStudy (mathmale):

We want to isolate t. We don't want t/2, but do want t alone. Therefore, as nincompoop asked, think about what you'd have to do to get rid of that 2 in the denominator on the left. Note that you must do the same operation to both sides of your inequality.

OpenStudy (mathmale):

Multiply both sides of the inequality \(\frac{ t }{ 2 }\le 4\) by (what??) to isolate t on the left side of this inequality.

OpenStudy (anonymous):

by 2

OpenStudy (mathmale):

Experiment. Don't answer until you're pretty sure you've got it. Yes, multiply both sides by 2. Would you please do that now? What does your inequality look like now?

OpenStudy (anonymous):

its 8

OpenStudy (anonymous):

its wat it looks like now

OpenStudy (mathmale):

@marrero, would you please be more specific? You start with an inequality; you multiply both sides by 2 to isolate t. You are still left with an inequality, but have not typed it. Please go back, multiply both sides of the inequality by 2, and then type out the entire resulting inequality.

OpenStudy (mathmale):

Excuse the reminder, Marrero, but the word "what" is spelled with an "h": "what," not as "wat."

OpenStudy (anonymous):

t is < or = to 8

OpenStudy (mathmale):

that's great. You have solved this inequality. Congratulations!

OpenStudy (anonymous):

ik i just like writeing it like that

OpenStudy (anonymous):

and thx

OpenStudy (mathmale):

It doesn't reflect well on you to misspell a common word like that. Of course you can do as you wish, but my recommendation would be that you always spell "what" as "what." I look forward to working with you again. what state are you in, by the way?

OpenStudy (anonymous):

Florida

OpenStudy (mathmale):

cool. Interesting state! See y ou again, later on! Take care.

OpenStudy (anonymous):

u 2